# An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks [Experiment, Analysis & Benchmark]

### v1

@@ 6.1.2. Workload Webpages
Next, we examine a workload that is drastically different from the DBLP workload. As shown in Section 5.1, Webpages has a significantly smaller number of queries in its workload compared to DBLP. Each string in the Webpages dataset is an HTML file for a webpage, making the average string length the longest among all the workloads. We run the workloads using the three methods and summarize the experimental results in Section 6.1.2.
We observed that the computational overhead of BEST is significantly reduced compared to DBLP due to the smaller query set. BEST achieves reasonable precision for performance improvement with a very small index size. However, since the query set is not representative, this small index may not benefit on unseen queries.
Looking at the first row of Section 6.1.2, by selecting only 3 to 4 n-grams for indexing with BEST($c=0.1$), the index achieves a precision of 0.138. By manually early stopping FREE($\max_{n}=2$,$c=0.02$) at 5 n-grams selected, it achieves comparable precision of 0.123, while incurring around 20$\times$ lower time overhead.
Table 4. Index cost and query performance on Webpages.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ GB $S_{I}$ MB $Prec$
5 BEST 422 8587 150 0.1 0.138
FREE 532 428 148 0.1 0.123
$\mathbf{1.7\times 10^{5}}$ BEST 422 8587 150 0.1 0.138
FREE 353 1184 159 357.9 0.163
$\mathbf{3.6\times 10^{5}}$ BEST 422 8587 150 0.1 0.138
FREE 240 5965 240 2007.4 0.301
$\mathbf{1.5\times 10^{6}}$ BEST 422 8587 150 0.1 0.138
FREE 239 7480 233 524.9 0.302
To attain a similar precision without early stopping, FREE($\max_{n}=2$,$c=0.5$) requires an index with more than 10000 times the number of keys and more than 3000 times the index size at $K=1.7\times 10^{5}$ The difference in the number of keys results from the different sources of n-grams considered during selection for the two methods. BEST chooses index keys from the intersection of n-grams in both the query set and the dataset, and a smaller number of queries means a smaller number of potential n-grams, effectively constraining the index size and computation overhead. On the other hand, FREE derives its index keys from the dataset, which contains many more distinct n-grams, leading to the selection of a large set of n-grams that meet the requirements defined by the specific configuration. Looking at the precision, we see that FREE, with a 0.025 higher precision than BEST, achieves a 1.2×1.2$\times$1.2 × overall query time speedup, reducing the time from 421.6 to 353.0 seconds.
When the number of keys allowed $K$ increases from $1.7\times 10^{5}$ to $3.6\times 10^{5}$ , the precision of the index by FREE($\max_{n}=2$,$c=0.7$) increases significantly from 0.163 to 0.301. The index construction time also increases significantly. The query time of the best index built by FREE for this $K=3.6\times 10^{6}$ is 1.47×1.47$\times$1.47 × faster and 1.75×1.75$\times$1.75 × faster than the index built with the three n-grams selected by BEST. Compared to the DBLP workload, the precision increase in Webpages corresponds to a more significant workload runtime decrease from 353 seconds to 240 seconds. In this workload, each webpage is long, and the cost of regex evaluation on one string is high. Therefore, eliminating more strings with the index significantly improves the performance. These advantages might not justify the large index size compared to BEST for this specific workload. However, it does make the index more robust to unseen queries on the same dataset. LPMS did not finish for this dataset within the set time frame due to the large average number of characters per line of the dataset. Unlike BEST where the data structures for intermediate calculation are adjacency lists due to the sparseness Assumption Assumption BEST-4, LPMS builds matrices of size $|Q|\times|G|$ as input to integer program solver, where the actual computational and space overhead is too large.
@@ 6.1.3. Workload Prosite
Prosite has the largest query size to number of records in the dataset ratio among the real-world workloads. The strings in the dataset have a mean length of 416 characters. Two distinguishing characteristics of Prosite are: 1) it has the smallest alphabet size of 22, and 2) it has very short literal components in its the regex queries.
Table 5. Index cost and query performance on Prosite.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
50 BEST 141.5 400 0.931 9.4 0.00826
FREE 154.0 2 0.227 1.5 0.00651
LPMS 151.8 14 0.263 26.2 0.00708
100 BEST 139.7 856 0.967 14.9 0.0089
FREE 153.6 2 0.242 5.7 0.00652
LPMS 151.8 14 0.263 26.2 0.00708
500 BEST 139.7 856 0.967 14.9 0.00890
FREE 150.8 3 0.595 141.4 0.00687
LPMS 151.8 14 0.263 26.2 0.00708
@@
Section 6.1.3 shows the results. BEST performs the best for this workload, generating index with highest precision while using a small number of keys of 364. The small alphabet size and small literal size per query make the number of possible n-grams considered in each calculation iteration small, and thus drastically reduce the index construction runtime overhead, which is the biggest advantage of BEST in other workloads.
@@
When indexing with 50 keys, FREE($\max_{n}=2$,$c=0.15$) generates an index with a higher precision of 0.00651, while taking the least computational overhead and smallest index size. LPMS-D, taking slightly longer to select the n-grams and generate the index, achieves a precision of 0.00708 within 14 seconds. It achieves similar query performance compared to the index by FREE, although the index size is significantly larger, as it selects more frequent n-grams as keys. BEST($c=0.7$) achieves the highest precision, 0.00826, among the three methods. Its query time is also more than 10 seconds lower than the other two methods. However, it requires 29×29$\times$29 × the index construction time compared to LPMS and 200×200$\times$200 × compared to FREE. It also consumes significantly more compute memory, nearly 3×3$\times$3 × higher than that of FREE and LPMS when $K=50$ and $K=100$. Overall index precisions are low for Prosite workload, as the average length of literals is very short in the regexes.
When $K=100$, the index by FREE($\max_{n}=2$,$c=0.2$) has almost identical precision with the 50-key index constructed by the same method. The workload running time for LPMS-D is slightly faster than FREE, with a similar compute memory. Index by BEST($c=0.7$) has a higher precision of 0.0089, resulting in a more than 10 seconds workload performance improvement over other methods. BEST takes 856 seconds to construct the index, much longer than other methods, but having smaller index size than LPMS. As the key size constraint loosens to $K=500$, FREE selects an index which only marginally decreases its overall workload running time from 153.6 seconds to 150.8 seconds, while using 395 keys (330 more keys than when $K=100$). The size of the index increases to 24.6×24.6$\times$24.6 × compared to the index built by FREE when $K=100$, and is much larger than the index sizes of other methods. However, its precision is still less than that of BEST and LPMS with a much smaller number of keys.
The results align with the characteristics of the methods: BEST trades off n-gram selection time for higher precision n-grams tailored to the workload, while LPMS approximates the results of BEST by trading off some precision for lower index construction overhead. When comparing the index construction memory and time overhead to the results for the same $K$ in DBLP, we observe that for Prosite, BEST has significantly smaller overhead that is not proportional to the dataset size difference. In Prosite, the possible literal set is small due to the limited alphabet size and short query literal sizes. Consequently, the computation time and space are significantly reduced in each step when BEST needs to count the number of lines for each n-gram.
@@
For workloads where query literals are short and/or the alphabet size is small, BEST can select an n-gram set with high filtering precision while incurring reasonable computational and storage overhead. This advantage arises from the decreased number of potential n-grams to consider. FREE does not benefit from this advantage since it only looks at n-grams in the dataset. LPMS benefits less from the characteristics of the workload due to the fixed overhead for constructing the integer program.
@@ 6.1.4. Workload US-Acc
Table 6. Index cost and query performance on US-Acc.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
30 BEST 1.451 63 2.204 84.80 0.510
FREE 4.940 29 0.969 0.02 0.032
LPMS 1.353 235 2.876 60.85 0.572
1000 BEST 1.451 63 2.204 84.80 0.510
FREE 1.708 31 3.235 1003.99 0.460
LPMS 1.353 235 2.876 60.85 0.572
5000 BEST 1.451 63 2.204 84.80 0.510
FREE 1.297 46 3.199 1070.97 0.811
LPMS 1.353 235 2.876 60.85 0.572
Section 6.1.4 shows the results. LPMS-D performs the best for the US-Acc workload, achieving a precision of 0.572 using only 12 keys. Although it takes 3.7×3.7$\times$3.7 × the index computation time compared to BEST($c=0.7$) and slightly more computation space overhead, LPMS uses 0.4×0.4$\times$0.4 × less index space to achieve this higher precision. When $K$ increases to 1000, FREE generates an index with 837 keys and a precision of 0.46, which is much lower than BEST and LPMS with fewer than 30 keys. This is because FREE selects n-grams based solely on their selectivity in the dataset, omitting those n-grams that are frequent in both the regex query and the dataset. In fact, the configuration of FREE that generates the highest precision with fewer than 1000 keys uses a selectivity threshold $c=0.7$, the largest across all of our experiments.
@@
For workloads where the query literals are very short and common in the dataset, LPMS can quickly locate a small set of n-grams with high precision. Both BEST and FREE remove n-grams that are frequent in the dataset, as they both adhere to Assumption Assumption FREE-1 that does not hold for the dataset.
@@
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
50 FREE 3457 1342 76.7 6643 0.000945
LPMS 1421 4569 147.5 4474 0.009855
5000 FREE 1473 1483 74.0 6643 0.002698
LPMS 1421 4569 147.5 4474 0.009855
@@ 6.1.5. Workload SQL-Srvr
The SQL-Srvr workload has dataset characteristics similar to the US-Acc workload, as it is generated from several formatted strings—in this case, system log reporting strings. We summarize the results in Section 6.1.4. The size of both the query set and dataset are much larger than those in US-Acc, and therefore, BEST failed to finish within the set time-frame.
LPMS performs better than FREE for similar reasons: it considers the query set without overlooking common n-grams. When we limit the number of n-grams to 50, LPMS-D selects an n-gram set with a precision of $9.855\times 10^{-3}$, which is 10×10$\times$10 × higher than the index generated by FREE($\max_{n}=2$,$c=0.7$) with the same number of keys. This is because FREE blindly selects selective n-grams without considering the queries. Given the variability in the values of the formatted system logs (e.g., VM IDs, cluster names), FREE may index short n-grams from them, resulting in an index that hardly benefits the actual query set. Consequently, the query time for LPMS is 2.43×2.43$\times$2.43 × faster. Although LPMS takes 3.4×3.4$\times$3.4 × longer and requires 1.9×1.9$\times$1.9 × more construction space and time overhead, the index built by LPMS uses 1.48×1.48$\times$1.48 × less space. Considering that LPMS and FREE use the same inverted index implementation and they have the same number of keys (50), the only difference in their index sizes comes from the posting lists lengths, which are the IDs of the data records. This suggests that LPMS selects n-grams that are more selective in the dataset. Combined with its higher precision and faster query time, LPMS selects n-grams that are both more selective in the dataset and more common in the query set, making them more beneficial. The results remain consistent until the upper limit of the number of n-grams grows to 5000. At this point, FREE($\max_{n}=2$,$c=0.03$) selects an n-gram set of size 4643, increasing the precision to $2.68\times 10^{-3}$. This is achieved with only a slight increase in index construction overhead and nearly identical index size compared to indexing with 50 n-grams selected by LPMS. However, the precision and query time are still worse than those of LPMS with 50 n-grams.
@@
Table 8. Index cost and query performance on Synthetic.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
20 BEST 0.829 3.398 0.042 0.091 0.2084
FREE 0.190 0.027 0.016 0.446 0.2672
LPMS 1.211 3.720 0.053 0.010 0.1479
100 BEST 0.444 31.659 0.043 0.452 0.4793
FREE 0.190 0.027 0.016 0.446 0.2672
LPMS 1.063 3.791 0.052 0.016 0.1757
300 BEST 0.061 90.661 0.043 1.125 0.6451
FREE 0.064 0.032 0.015 1.122 0.6453
LPMS 1.063 3.791 0.052 0.016 0.1757
6.1.6. Robustness Test
For this test, we use the synthetic workload with different sets of regexes for index construction and query matching to test the robustness of each method for unseen queries. We summarize the result in Section 6.1.5.
For $K=20$, FREE($\max_{n}=2$, $c=0.7$) builds an index with 16 keys that achieves the highest precision of 0.267. It also has the lowest index construction overhead. BEST($c=0.2$) achieves slightly lower precision of 0.208, but the query time is $\times4$ more than that of FREE. LPMS-D performs the worst with the lowest precision and highest query time. Note that BEST and LPMS also have significantly higher computational overhead compared to FREE.
As $K$ increases to 100, BEST($c=0.5$) achieves the highest precision. It’s index construction time also increases by $\approx$3 times to 31.7s. This is more than 1000×1000$\times$1000 × longer than FREE. Note that the query times for BEST are not the lowest among the three methods, although it has the highest precision. This is due to the non-uniformness in the data record lengths, which breaks Assumption Assumption BEST-2.
As we increase to $K=300$, FREE($max_{n}=2$, $c=0.12$) achieves the highest precision of 0.6453. BEST($c=0.2$) achieves similar precision and similar overall query time compared to FREE, but with a much larger index construction time which is 2833×2833$\times$2833 × slower than FREE. In all cases, LPMS performs unfavorable to the other two methods, since the n-gram selection strategy of LPMS is solely based on the benefit calculation using the query set and the dataset, without any heuristic like the pruning step of BEST.

### v2

@@ 6.1.2. Workload Webpages
We evaluate DBLP, a workload with significantly fewer queries than DBLP, using HTML files from Webpages, which has the longest average string length among all workloads. Results are summarized in Section 6.1.1, with a baseline (no index) query time of 437s.
At $K=5$, BEST($c=0.1$) selects just four n-grams and achieves 0.138 precision. A manually early-stopped FREE($\max_{n}=2$, $c=0.02$) with five n-grams yields a similar 0.123 precision. To match that precision without early stopping, FREE($\max_{n}=2$, $c=0.5$) requires $1.7\times 10^{5}$ keys. This difference stems from selection strategies: BEST chooses n-grams from the query–dataset intersection, limiting candidates when queries are few. In contrast, FREE selects from the dataset alone, resulting in a much larger candidate pool. Despite requiring more keys, FREE achieves 0.025 higher precision than BEST and delivers a 1.2× speedup in query time—reducing it from 421.6s to 353.0s as shown in Section 6.1.1.
When the number of keys allowed $K$ increases from $1.7\times 10^{5}$ to $3.6\times 10^{5}$ , the precision of the index by FREE($\max_{n}=2$,$c=0.7$) increases significantly from 0.163 to 0.301. At $K=3.6\times 10^{5}$, FREE delivers a query time that is 1.47× faster than its own earlier version and 1.75× faster than BEST. Compared to the overall precision and query time trend for DBLP workload in Figure 2, the precision increase of FREE in Webpages corresponds to a more significant workload runtime decrease from 353 seconds to 240 seconds. Since each string in Webpages is a long HTML page, filtering more strings upfront yields greater performance benefits. While FREE’s larger index may not always justify its size relative to BEST, it offers better robustness to unseen queries on the same dataset. Due to lack of space, we defer the reader to Appendix B of the full paper for detailed experiments. VGGraph and LPMS did not finish for this dataset within the set time frame due to the large average length per string of the dataset.
@@ 6.1.3. Workload Prosite
Prosite has the largest query size to number of records in the dataset ratio among the real-world workloads. The strings in the dataset have a mean length of 416 characters. Two distinguishing characteristics of Prosite are: 1) it has the smallest alphabet size of 22, and 2) it has very short literal components in its the regex queries. The short literal components gives little opportunity for n-gram index improvement, and all methods created indices with low precision and performance improvement. In fact, the baseline workload runtime is 154 seconds.
(a) Precision
(b) Query Time $T_{Q}$ (s)
Figure 3. Prosite.
@@
BEST is the most performant for this workload, when $K$ is small, generating index with highest precision while using only 364 keys. The small alphabet size and small literal size per query make the number of possible n-grams considered in each calculation iteration small, and thus drastically reduce the index construction runtime overhead, which is the biggest advantage of BEST in other workloads. Similarly, VGGraph which adapts a similar cover strategy, achieves the best precision as $K$ gets larger.
@@
When indexing with 50 keys, FREE($\max_{n}=2$,$c=0.15$) generates an index with a higher precision of 0.00651, while taking the least computational overhead and smallest index size. LPMS-D, taking slightly longer to select the n-grams and generate the index, achieves a precision of 0.00708 within 14 seconds. BEST($c=0.7$) achieves the highest precision, 0.00826, among the three methods. Its query time is also more than 10 seconds lower than the other two methods. Overall index precisions are low for Prosite workload, as the average length of literals is very short in the regexes.
When $K=100$, FREE, LPMS, and Trigrams does not lead to much improvement. Index by BEST($c=0.7$) has a higher precision of 0.0089, resulting in a more than 10 seconds workload performance improvement over other methods. VGGraph ($max_{n}=6$, $c=0.7$) also increases slightly to 0.0732. The precision increments for all methods does not corresponds to significant performance improvement as shown in the bars in Figure 3(b). As the key size constraint relaxes to $K=500$, selected n-grams of VGGraph ($max_{n}=2$, $c=0.7$) achieves the highest precision of 0.0987, and the significant precision improvement from the previous $K$ led to a query time reduction to 128.5 seconds, $1.13\times$ query time reduction from the baseline. Since the literal lengths are small in this workload, the effectiveness of each n-gram is much more important for effective filtering. Therefore, n-gram sets selected by dataset-based methods, FREE and Trigrams, are not effective in filtering.
@@
For workloads where query literals are short and/or the alphabet size is small, BEST can select an n-gram set with high filtering precision while incurring reasonable computational and storage overhead. This advantage arises from the smaller n-gram candidate set. VGGraph outperform other methods in terms of precision as number of n-grams scales up. FREE does not benefit from this advantage since it only looks at n-grams in the dataset. LPMS benefits less from the characteristics of the workload due to the fixed overhead for constructing the integer program.
(a) Precision
(b) Query Time $T_{Q}$ (s)
Figure 4. US-Acc.
@@ 6.1.4. Workload US-Acc
Figure 4 shows the results. The methods that uses both query set and dataset as n-gram sources generate better query performance than the baseline query time of 4.8 seconds. LPMS-D performs the best for the US-Acc workload, achieving a precision of 0.572 using only 12 keys. When $K$ increases to 1000, FREE generates an index with 837 keys and a precision of 0.46, which is much lower than BEST and LPMS with fewer than 30 keys. This is because FREE selects n-grams based solely on their selectivity in the dataset, omitting those n-grams that are frequent in both the regex query and the dataset. When $K$ increases to 5000, VGGraph ($max_{n}=3$, $c=0.7$) generates a index with 4247 keys and achieves an query time of 1.51 seconds that is comparable to other faster methods.
@@
For workloads where the query literals are very short and common in the dataset, LPMS can quickly locate a small set of n-grams with high precision. VGGraph, BEST, and FREE remove n-grams that are frequent in the dataset, as they both adhere to Assumption Assumption FREE-1 that does not hold for the dataset.
@@
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
[…]
@@ 6.1.5. Workload SQL-Srvr
The SQL-Srvr workload, like US-Acc, consists of structured strings—specifically, system log reports—but with a much larger dataset and query set. BEST was unable to complete within the time limit. Results are summarized in Section 6.1.4.
[…]
@@
6.1.6. Workload Enron
The Enron workload contains long data strings in the dataset like Webpages, yet it has relatively small dataset size and larger query size. The lengths of literal components among queries are generally short. Although each data string is long, the small dataset size and the short literal components make the computation of majority of the methods within our set time limit. We run BEST, FREE, LPMS, VGGraph, and Trigrams on this dataset and summarize the results in Section 6.1.6. Without indexing, Enron runs in 126 seconds as a baseline.
BEST ($c=0.7$) achieves highest precision for $K=10$. With a precision of 0.148, around $1.4\times$ higher than other methods, BEST uses $1.2\times$ less time for the workload querying. As $K$ increases to 100, precision of BEST increases by $2.2\times$ to 0.321, and the query time decreases by $1.5\times$. LPMS-D, having similar strategy as BEST, also achieves precision incremental as $K$ increases from 10 to 100 with significantly reduced precision. Enron has one of the highest regex match rate, where without filtering, the true positive rate per regex is already around 0.105, suggesting that VGGraph and Trigrams does not perform effective filtering.
When number of keys selected for the index continue to increase, the query performance and precision fails to increase proportionally. This is due to the fact that literal components in the queries are fairly common in the data strings, getting filtered out due to the selectivity threshold earlier on.
@@ 6.1.6. Workload Enron
BEST demonstrates exceptional performance in Enron, efficiently leveraging its optimization-based approach with smaller query sets and longer literals. VGGraph shows substantial performance enhancement at moderate index sizes. LPMS offers balanced precision and performance, while FREE and Trigrams indexing lag behind due to their less targeted selection methods.
Table 5. Index cost and query performance on Enron.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
[…]
LPMS 124 79 3 16.3 0.107
[…]
FREE 123 35 3 5.1 0.109
LPMS 101 2104 13 256.2 0.149
[…]
6.1.7. Workload ADX
The ADX workload has the largest dataset and a very small query set. Trigrams method would not generate useful index if tight constraint is posed the number of n-grams selected, and will generate index too large to fit into memory if no constraint imposed. VGGraph did not finish for all configuration on this workload. This workload runs in 1663 seconds without indexing. We run FREE, BEST, and LPMS on ADX workload and summarize the result in Section 6.1.7.
Due to the small query set size, and large and diverse data strings in the dataset, methods considering both query and dataset, BEST and LPMS achieves better precision and performance improvement for all key constraints $K$. Overall, LPMS-D achieves best precision of 0.328 and query performance of 221 seconds, $7.52\times$ faster than the baseline. BEST ($c=0.1$) also achieves $5.3\times$ to $6.5\times$ performance improvement with 10 to 25 keys.
{customframe}
Insights
For ADX, a workload with simple patterns and very large dataset size, LPMS stands out by effectively balancing precision and computational overhead. BEST becomes computationally prohibitive due to its complexity, and Trigrams indexing is impractical at this scale.
Table 6. Index cost and query performance on ADX.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
[…]
LPMS 373 2184 104 2799.0 0.113
25 BEST 254 28912 219 650.6 0.216
[…]
6.2. Index Construction Overhead
In this section, we aim to answer question Q.2 with a detailed analysis of the index construction overhead for FREE, BEST, LPMS, VGGraph, and Trigrams, based on their time and space usage observed during index construction.
FREE demonstrates a significantly lower index construction time across various workloads due to its straightforward selectivity-based strategy and iterative prefix-free minimal n-gram selection approach. Specifically, for the Webpages workload, as shown in Section 6.1.1, FREE constructed an index with $1.7\times 10^{5}$ keys in merely 1184 seconds compared to BEST’s significantly longer 8587 seconds. FREE’s runtime space overhead remains relatively constant and low across workloads, generally proportional to the dataset size due to the prefix-free minimal set construction method.
BEST has a high index construction overhead because it performs an exhaustive search to approximate optimal n-gram set through a utility-based set-cover strategy. Specifically, in the DBLP workload, BEST required 8762 seconds to construct an index of 363 keys, dramatically more than the sub-second construction time achieved by FREE for a comparable number of keys. The high overhead results from the auxiliary data structures mentioned in Section 4.2.2, required to maintain for time-efficient computation. In workloads with larger query sets or longer literals, such as Webpages in Section 6.1.1, BEST’s computation overhead increases and potentially exceeding practical time limits.
LPMS has moderate to high index construction times, often higher than FREE but lower than BEST. For instance, in the SQL-Srvr workload results in Section 6.1.4, LPMS required 4569 seconds for constructing an index with 50 keys, significantly exceeding FREE’s 1342 seconds for an similar number of keys but less than BEST. LPMS’s overhead results from its formulation and solving of the linear program.
VGGraph has an index construction overhead generally lower than BEST but higher than FREE. The overhead is attributed to its iterative evaluation of cost-efficiency ratios. For the SQL-Srvr workload in Section 6.1.4, VGGraph constructed indices in 45 seconds, considerably lower than BEST and LPMS that runs more than 2000 seconds but slightly higher than FREE (30s). Its overhead scales as the dataset and query complexity increase. Trigrams has minimal index construction overhead when limited to a small number of keys, but this overhead grows rapidly as more trigrams are indexed—especially without key constraints. Its cost is primarily driven by dataset size rather than query complexity or optimization.
(a) Precision
(b) Index Time $T_{I}$ (in seconds)
Figure 5. Synthetic workload with $|D|=50,000$ and average query selectivity 0.02.
6.3. Case Study: Workload Characteristics
In this section, we perform an ablation study with Synthetic workload to evaluate how different regex indexing methods behave under varying query set sizes and query selectivity. For each configuration, we constrain the maximum number of n-grams ($K$) to 5000 and measured both precision and query runtime ($T_{Q}$). BEST failed to finish within the time limit. However, in the experimental result, varying regex selectivity does not result in a common trend on the indices built with different methods in workload performance or index construction. We will focus on the analysis of varying query size with fixed dataset size and regex query selectivity.
Figure 5(a) illustrates that as the query set size increases from 1000 to 5000, FREE consistently maintained high precision. This can be attributed to the small alphabet size of the workload. LPMS and VGGraph also showed stable but lower precisions, with LPMS ranging around 0.95 and VGGraph at approximately 0.96. All three methods achieves similarly high precision. The precision of all methods slightly improve as query set size increased, as a larger query set provides more opportunities for n-gram filtering. Regarding index construction time ($T_{I}$), in Figure 5(b), FREE and VGGraph exhibited significantly lower overhead, maintaining fast index build times which is under 1 second across all query set sizes. LPMS required considerably more time as the query set size increased. LPMS is sensitive to query set growth, requires over 90 seconds at larger query sets. This matches our observation from Section 6.2, where LPMS and BEST incurred much larger index construction overhead.